OpenStudy (anonymous):

check answer please 5x^2-5x+1 solve using the quadratic formula

OpenStudy (kainui):

What answer did you get?

OpenStudy (anonymous):

\[x=-5\pm \frac{ \sqrt{-5^{2}-4(5)(1)} }{ 2(5) }\]

OpenStudy (anonymous):

\[x=-5\pm \frac{ \sqrt{45} }{ 10 }\]

OpenStudy (anonymous):

but in my answer choices i dont have this do i simplify the 45?

OpenStudy (noelgreco):

The negative sign is incorrect.

OpenStudy (anonymous):

i know typo

OpenStudy (noelgreco):

\[\sqrt{45}=\sqrt{9}\times \sqrt{5}\]

OpenStudy (anonymous):

so my answer would be \[x=\frac{ 5\pm \sqrt{5} }{ 10 }\]

OpenStudy (noelgreco):

What happened to the 3 in front of the radical?

OpenStudy (anonymous):

what?

OpenStudy (noelgreco):

\[\sqrt{9}=3\]

OpenStudy (anonymous):

ok im conffused

OpenStudy (kainui):

After you square root 9, 3 vanishes like a ghost for no reason. He's like, "I did my business now I'm GONE!"

OpenStudy (noelgreco):

\[\sqrt{45}=\sqrt{9\times5}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\]

OpenStudy (anonymous):

what would my final answer be

OpenStudy (noelgreco):

Just put a 3 to left of the radical (square root sign) in your last answer.

OpenStudy (anonymous):

ok my answer choices dont have a 3 in front of the square root sign

OpenStudy (anonymous):

answer choices \[x=\frac{ 5\pm \sqrt{5} }{ 10},x=\frac{ -8\pm \sqrt{28} }{ 6 },x=\frac{5\pm \sqrt{45} }{ 20 }\]

OpenStudy (anonymous):

this is how my question looks like

OpenStudy (anonymous):

Which two values of x roots are polynomial below 5x^2+5x-1